An ice cube of mass 65 g is taken from a freezer at -15o C and dropped into 125 cm3 of water at 85o C. What will be the final temperature of the mixture? The specific heat of ice is 0.5 cal/g.o C and latent heat of fusion of ice is 80 cal/g.


1
Expert's answer
2022-03-02T03:10:09-0500

m1s1(T1-T)= m2s2(0-T2) + m2L + m2s2 (T-T2)

125×1×(85-T)= 65×0.5(0- -15) + 65×80 + 65×1×(T-0)

10625-125T = 487.5+5200+65T

190T = 4937.5

T= 25.987°C

Final Temperature is;

= 25.987

= 26°C